OpenStudy (anonymous):

Solve. 3/2x – 4 = 16 (1 point) x = 20 x = 13* x = 15 30 WILL GIVE MEDAL FOR CORRECT ANSWER!!!!

OpenStudy (anonymous):

3/2x - 4 = 16 Add 4 to both sides: 3/2x = 20 Divide: x = 3.3

OpenStudy (anonymous):

Are you sure you wrote the equation down right?

OpenStudy (nirmalnema):

is it 3/2 x - 4 = 16